OpenStudy (anonymous):

Use the following functions, to find the three answers. f(x) = x + 5 g(x) = x2 -4x + 5 h(x) = 3x2 - 2x + 5 Find (f+g)(x) Find (f*h)(x) Find f [g(2)]

OpenStudy (anonymous):

just plug in :)

OpenStudy (jdoe0001):

yeah (f + g)(x) => f(x) + g(x)

OpenStudy (anonymous):

Distribute (f+g)(x)=f(x)+g(x)

OpenStudy (anonymous):

(f*h)(x) f(x)*h(x) you dont really distribute but that is what it really means. Shorthand i think.

OpenStudy (anonymous):

Oh...I get it now thank you, my math book doesn't really explain it at all.

OpenStudy (anonymous):

so the first one. If f(x)=x + 5 and g(x) = x2 -4x + 5 then f(x) + g(x) ... plug in ^ ^ (x + 5 ) + (x2 -4x + 5) now simply add like terms. SInce it is addition you can get rid of the perenthesis.

OpenStudy (anonymous):

Thank you that explains it really good

OpenStudy (anonymous):

Find (f*h)(x) that one is almost the same f(x)*h(x) so multiply f(x) = x + 5 h(x) = 3x2 - 2x + 5 plug in (x + 5)*(3x2 - 2x + 5) here the perenthesis do matter so distribute the x+5 :) Let me know if you need help with that

OpenStudy (anonymous):

THis one is a lil more complex but still easy Find f [g(2)] you got g(2) inside of f(x) right? g(x)= x2 -4x + 5 However you probably know when you are given a number you evaluate g(2)= (2)^2 -4(2) + 5= 4-8+5= -4+5=1 so g(2)=1 that means f [g(2)] and if g(2)=1 then f(1) now you simply plug in f(x) = x + 5 is f(1) = (1) + 5 =6 so f [g(2)]=6 if i didn't make any careless mistakes. THaat is called a composite function you will also see it written as (f o g) (2) = f [g(2)] the o is the composite operator i think and i think that is the correct notation i dont really like it but i think that is the right alternative notation you will see. Remember always do the inner most part first :D

OpenStudy (anonymous):

Thanks this helps me a lot thanks